OpenStudy (anonymous):

Divide -2x^3+21x^2-20x-12 by (x-3)

OpenStudy (opcode):

\[p(x) = -2x^{3}+21x^{2}-20x-12\] \[q(x) = x-3\] Use the synthetic division table: \[\begin{matrix}\begin{array}{r} 3 \\ ~ \end{array} & \underline { \begin{array}{rrrr} -2&21&-20&-12\\ ~~~~&-6&45&75 \end{array} } \\ ~ & \begin{array}{rrrr} \color{blue}{ -2 }&\color{blue}{ 15 }&\color{blue}{ 25 }&\color{red}{ 63 } \end{array}\end{matrix}\] \[\frac{ -2x^{3}+21x^{2}-20x-12 }{ x-3 } = \color{blue}{ -2 }x^{2}~\color{blue}{ + }~\color{blue}{ 15 }x~\color{blue}{ + }~\color{blue}{ 25 } ~+~ \frac{ \color{red}{ 63 } }{ x-3 }\]

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (opcode):

Do you understand?

OpenStudy (anonymous):

Somewhat, yeah.

OpenStudy (opcode):

Is there anything I can do to help clarify?

OpenStudy (anonymous):

I'm not really able to do normal division, even. This is about as good as I'm going to get on it.

OpenStudy (anonymous):

Thank you though.

OpenStudy (opcode):

The Khan Academy video explain really well on what I did, I highly recommend you watch it.

OpenStudy (anonymous):

I will, thanks!

OpenStudy (opcode):

Good luck!